# Empirical universality and non-universality of local dynamics in the Sherrington-Kirkpatrick model

### v1

@@ Abstract
Several recent works have aimed to design algorithms for optimizing the Hamiltonians of spin glass models from statistical physics. While Montanari (2018) eventually gave a sophisticated message-passing algorithm to do this nearly optimally for the Sherrington-Kirkpatrick (SK) model, Parisi (2003) observed earlier that a simple yet unusual algorithm seems to perform just as well: perform local reluctant search, repeatedly making the local adjustment improving the objective function by the smallest possible amount. This is in contrast to the more intuitive local greedy search that repeatedly makes the local adjustment improving the objective by the largest possible amount. We study empirically how the performance of these algorithms depends on the distribution of entries of the coupling matrix in the SK model. We find evidence that, while the runtime of greedy search enjoys universality over a broad range of distributions, the runtime of reluctant search surprisingly is not universal, sometimes depending quite sensitively on the entry distribution. We propose that one mechanism leading to this non-universality is a change in the behavior of reluctant search when the couplings have discrete support on an evenly-spaced grid, and give experimental results supporting this proposal and investigating other properties of a distribution that might affect the performance of reluctant search.
@@ 1.2 Local and Non-Local Algorithms
We will focus on a family of even simpler local algorithms, which we describe below. Though they may seem naive, we will see that even among these algorithms there are interesting and surprising behaviors.
@@
We adopt the convention that, if there are no feasible $\sigma^{\prime}$ for the $\operatorname*{arg\,max}$, then $F_{\mathrm{rel}}(\sigma,J)$ returns $\sigma$. A reasonable intuition for why this is a good idea is that it is advantageous to “slow down” greedy descent on a complicated optimization landscape, in order to explore more of the energy landscape during the optimization process. Empirical studies indeed have verified that the reluctant algorithm consistently finds lower-energy configurations than the greedy algorithm, and non-rigorous analysis using the methods of statistical physics in fact suggests that the reluctant algorithm might also with high probability find configurations with energy arbitrarily close to the ground state energy [Par03, BCDEG03, Hor07, EBKZ24]. This would mean that the very simple reluctant local algorithm indeed matches the performance of the far more complicated algorithm of [Mon18].
@@ 2.1 Estimating Runtime Scaling Laws
$\mu$ $\widehat{\beta}(\mu,0)$ $\widehat{\beta}(\mu,\infty)$ $\Delta(\mu)$ $\mathrm{supp}(\mu)$ $\mathbb{P}(J_{ij}=0)$
Moments
Matching
$\mathcal{N}(0,1)$
@@ 2.2 Choice of Distributions
3 Numerical Results: Scaling Exponents
@@ 3.5 Sparsity
We give the results in Figure 8. These clearly demonstrate that the scaling exponent $\widehat{\beta}(\mathrm{Sparsify}(\mu,p),0)$ of greedy dynamics does not depend on $p$, consistent with Conjecture 1.4. This is less clear for small values of $p$, but in this case as the coupling matrix becomes very sparse we suspect that various high-dimensional concentration of measure phenomena might not yet “kick in.” We leave a more careful investigation of this matter with larger-scale experiments to future work.
@@
4 Distributions of Energy Increments
Finally, we present some calculations suggesting a possible explanation for why the discrepancy of a measure is so closely related to the behavior of reluctant dynamics on coupling matrices with that distribution of entries. Namely, we consider the distribution of the size of the first step taken by greedy and reluctant dynamics from a uniformly random initialization.
@@ Acknowledgments
Thanks to Ilias Zadik and Florent Krzakala for bringing reluctant dynamics to our attention, and to Cristopher Moore for helpful discussions.

### v2

@@ Abstract
Several recent works have aimed to design algorithms for optimizing the Hamiltonians of spin glass models from statistical physics. While Montanari (2018) eventually gave a sophisticated message-passing algorithm to do this nearly optimally for the Sherrington-Kirkpatrick (SK) model, the recent work of Erba, Behrens, Krzakala, and Zdeborová (2024) also observed that a simple yet unusual algorithm first proposed by Parisi (2003) seems to perform just as well: perform local reluctant search, repeatedly making the local adjustment improving the objective function by the smallest possible amount. This is in contrast to the more intuitive local greedy search that repeatedly makes the local adjustment improving the objective by the largest possible amount. We study empirically how the performance of these algorithms depends on the distribution of entries of the coupling matrix in the SK model. We find evidence that, while the runtime of greedy search enjoys universality over a broad range of distributions, the runtime of reluctant search surprisingly is not universal, sometimes depending quite sensitively on the entry distribution. We propose that one mechanism leading to this non-universality is a change in the behavior of reluctant search when the couplings have discrete support on an evenly-spaced grid, and give experimental results supporting this proposal and investigating other properties of a distribution that might affect the performance of reluctant search.
@@ 1.2 Local and Non-Local Algorithms
While we do not consider it here, also worth mentioning is the slightly different framework of sequential local algorithms studied by [DGPZ25] (one instance of which was also compared to the algorithms we do focus on by [Par03]). These algorithms repeatedly scan through the spins in a fixed order and update them according to local information around the spin under consideration (in particular, in that work, according to rules involving the local field acting on spin $i$, which is the quantity $(J\sigma)_{i}$). Though this is a quite restrictive structure to demand of an algorithm, [DGPZ25] show that such algorithms, unlike (at least with current tools) those we propose to study empirically below, are amenable to rigorous theoretical analysis.
We will focus instead on a family of even simpler local algorithms, which we describe next. Though they may seem naive, we will see that even among these algorithms there are interesting and surprising behaviors.
@@
We adopt the convention that, if there are no feasible $\sigma^{\prime}$ for the $\operatorname*{arg\,max}$, then $F_{\mathrm{rel}}(\sigma,J)$ returns $\sigma$. A reasonable intuition for why this is a good idea is that it is advantageous to “slow down” greedy descent on a complicated optimization landscape, in order for the trajectory of iterates to have the opportunity to explore more of the landscape. Empirical studies indeed have verified that the reluctant algorithm consistently finds lower-energy configurations than the greedy algorithm [Par03, BCDEG03, Hor07], and, while at first smaller experiments (in particular those carried out by [Par03]) suggested otherwise, the recent work of [EBKZ24] found that both larger experiments and non-rigorous analysis using the methods of statistical physics in fact indicate that the reluctant algorithm might with high probability find configurations with energy arbitrarily close to the ground state energy. This would mean that the very simple reluctant local algorithm indeed matches the performance of the far more complicated algorithm of [Mon18].
@@ 2.1 Estimating Runtime Scaling Laws
$\mu$ $\widehat{\beta}(\mu,0)$ $\widehat{\beta}(\mu,\infty)$ $\Delta(\mu)$ $\mathrm{supp}(\mu)$ $\mathbb{P}(J_{ij}=0)$ Moments Matching $\mathcal{N}(0,1)$
@@ 2.2 Choice of Distributions
3 Main Numerical Results: Runtime Scaling
@@ 3.5 Sparsity
We present the results in Figure 8. These clearly demonstrate that the scaling exponent $\widehat{\beta}(\mathrm{Sparsify}(\mu,p),0)$ of greedy dynamics does not depend on $p$, consistent with Conjecture 1.4. This is less clear for small values of $p$, but in this case as the coupling matrix becomes very sparse we suspect that various high-dimensional concentration of measure phenomena might not yet “kick in.” We leave a more careful investigation of this matter with larger-scale experiments to future work.
@@
4 Additional Results: Distributions of Energy Increments
We next present some calculations suggesting a possible explanation for why the discrepancy of a measure is so closely related to the behavior of reluctant dynamics on coupling matrices with that distribution of entries. Namely, we consider the distribution of the size of the first step taken by greedy and reluctant dynamics from a uniformly random initialization.
@@ 4.2 Sizes of First Steps
5 Additional Results: Final Energy Levels
It is reasonable to wonder whether not only the runtime but the actual performance of reluctant dynamics is non-universal in terms of the final energy achieved after the algorithm converges. This would be quite surprising, since the asymptotic ground state energy itself is known to be universal across broad classes of coupling distributions [CH04], and at least for Gaussian couplings reluctant dynamics are believed according to the most detailed numerical studies to (nearly) reach that ground state energy [EBKZ24]. Thus, if reluctant dynamics performed worse for other distributions, it would also not share the qualitative property of (nearly) reaching the ground state energy, which seems unlikely.
In Figure 10, we present numerical experiments on the same scale as those substantiating our other predictions for the energy reached by greedy and reluctant dynamics for various coupling distributions, as a function of the system size $N$. The results are inconclusive: the difference in final energy level between different distributions, particularly between distributions with zero discrepancy and those with positive discrepancy, is greater for reluctant dynamics than greedy dynamics, but also reluctant dynamics is still quite far from reaching the true asymptotic ground state energy for the system sizes we consider, and thus remains far from its $N\to\infty$ asymptotic behavior (at least per the predictions of [EBKZ24]). Thus, it seems considerably larger experiments would be required to understand whether the final energy reached by reluctant dynamics is indeed non-universal, and we leave such further numerical studies to future work.
Figure 10: Final energy level at convergence. We compare the energy level achieved at convergence by greedy and reluctant algorithms for both zero and positive-discrepancy distributions of couplings. Dashed lines in all plots indicate the theoretical asymptotic ground state energy as $N\to\infty$.
@@ Acknowledgments
Thanks to Reza Gheissari, Florent Krzakala, and Ilias Zadik for bringing reluctant dynamics to our attention, to Cristopher Moore for helpful discussions, and to the authors of [DGPZ25] for pointing out the related results in their paper as well as the predictions of [EBKZ24] concerning reluctant dynamics.